OpenStudy (migitmack):

Given the function f(x) = 5x - 9, find f(-3). I think the answer is -24

Parth (parthkohli):

Correct!

OpenStudy (anonymous):

f(x) = 5x - 9 f(-3) = 5(-3) - 9 f(-3) = -15 - 9 f(-3) = -24 Yup you got it

OpenStudy (migitmack):

ok thx